OpenStudy (anonymous):

Verify Work. Can someone check to see if my solution is correct. Retyping using equation editor.

OpenStudy (anonymous):

\[\lim_{x \rightarrow -2}\frac{ x+2 }{ x^3+8 } = \frac{ 1 }{ 12 }\] This is how, I broke it down. \[\lim_{x \rightarrow -2}\frac{ (x+2) }{ (x+2)(x^2-2x+4) }\] I am asking for help because the book says the answer is -(1/16)

OpenStudy (anonymous):

that is correct. The book has a typo in it.

OpenStudy (anonymous):

Thanks